# Recalibrating binary probabilistic classifiers

### v2

@@ 5 Example
  • The mean of the square root of the class 1 posterior probabilities as an expample involving a concave function of the posterior probabilities.
For the example we chose discrete source and target distributions of a feature (called score in the following) with values in an ordered set with 17 elements. These distributions may be interpreted as empirical distributions of samples with many ties or as the genuine distributions of discrete scores or ratings. For instance, the major credit rating agencies Standard & Poor’s, Moody’s and Fitch use rating scales with 17 to 19 different grades.
The source distribution is specified as follows: (1) The conditional feature distribution for class 0 is a binomial distribution with success probability 0.4, the conditional feature distribution for class 1 is a binomial distribution with success probability 0.55. The number of trials for both binomial distributions is 16 such that the support of the distribution is the set $\{0,1,\ldots,16\}$. (2) The class 1 prior probability is $p=0.01$.
The target distribution is incompletely specified as follows: (1) The unconditional feature distribution is a binomial distribution with number of trials 16 whose success probability is a Vasicek-distributed random variable (Section 2.2 of Meyer [13]) with mean 0.3 and correlation parameter 0.3. (2) The target class 1 prior probability is $q=0.05$.
Figure 1 shows [^3] Calculations were performed with R (R Core Team [17]). Details of some more involved calculations are described in Appendices 0.A.1 and 0.A.2. The R-scripts can be downloaded from https://www.researchgate.net/profile/Dirk-Tasche. the source and target unconditional score distributions. They were intentionally chosen to be quite different such that any distribution shift assumed for the recalibration must be significant.
Figure 1: Source and target score distributions for the example in Section 5.
@@
Figure 2: Class 1 posterior probabilities for the example in Section 5. The scale of the vertical axis is logarithmic. The dots for the probabilities have been connected with straight lines for better readability. ‘Source’ refers to the posterior probabilities in the source distribution without recalibration.
All target posterior probabilities are well above the source posterior probabilities. This does not come as a surprise given that the target class 1 prior probability of 5% is much higher than the source class 1 prior probability of 1%. Otherwise, there are two groups of target posterior probability curves: The curves based on capped scaling, label shift and FJS on the one hand, and the curves based on the five QMM methods described in Sections 4.4 and 4.5 on the other hand. The curves of the former group are rather steep compared to the curves of the latter group.
Table 1 displays three characteristics for the source distribution as well as for the eight different target distributions that result from the recalibration methods discussed in Section 4. Concluding from the entries in the column ‘mean(probs)’, only the recalibration method ‘label shift’ – introduced in Section 4.2 – is unreliable in so far as it does not achieve the required target class 1 prior probability $p=0.05$. With all seven other recalibration methods the target prior probability is matched.
Column ‘AUC’ of Table 1 is more varied than column ‘mean(probs)’. For the five QMM methods from Sections 4.4 and 4.5, AUC is essentially the same as the AUC of 0.802 under the source distribution. This is a consequence of the QMM design since one of the moment matching objectives is hitting the source AUC. Thanks to Proposition 1, one might expect also the ‘label shift’ AUC to match the source AUC. However, Proposition 1 is not applicable to the example of this section because by design the target feature distribution is not a mixture of the source class-conditional distributions. In any case, the high AUC values for methods ‘capped scaling’, ‘label shift’ and ‘FJS’ cause the greater slopes of their posterior probability curves in Figure 2 compared to the slopes of the QMM posterior probability curves.
Table 1: Results for the recalibration methods presented in Section 4. The row ‘Source’ shows the values for the source distribution without recalibration. The numbers in all other rows refer to the target distribution. ‘mean(probs)’ is the class 1 prior probability calculated as mean of the recalibrated posterior probabilities. ‘AUC’ is the area under the ROC curve implied by the recalibrated posterior probabilities. ‘mean(sqrt(probs))’ is the mean of the square root of the recalibrated posterior probabilities.
@@
As an example for the application of a concave function to the target posterior probabilities, we chose the function $C(u)=\sqrt{u}$, $u\in[0,1]$, and the related expected value $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target distribution. Note that (2) implies $0.05=q\leq E_{Q}\bigl{[}\sqrt{\eta_{Q}(X)}\bigr{]}\leq\sqrt{q}\approx 0.2236$.
@@ 6 Conclusions
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. This paper presents analyses and methods for recalibration from a distribution shift perspective. It has turned out that distribution shift assumptions linked to the performance in terms of area under the curve (AUC) of a probabilistic classifier under the source distribution are useful for the design of meaningful recalibration methods. Two new methods called parametric CSPD (covariate shift with posterior drift) and ROC-based QMM (quasi moment matching) have been proposed and have been tested together with some other methods in an example setting. The outcomes of this testing exercise suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
{credits}
6.0.1 Acknowledgements
The author would like to thank two anonymous reviewers for their useful comments and suggestions.
Appendix 0.A Appendix
0.A.1 The AUC of discrete auto-calibrated probabilitistic classifiers
@@
How to calculate implied AUC for an auto-calibrated [^4] The probabilistic classifier $S$ is auto-calibrated if for all $s$ in the range of $S$ it holds that $\mu[Y=1\,|\,S=s]=s$ (Tsyplakov [26], Section 2.2). probabilistic classifier $S$ under a joint distribution $\mu(S,Y)$ of $S$ and $Y\in\mathcal{Y}=\{0,1\}$? We assume here that the distribution of $S$ under distribution $\mu$ is discrete, i.e. it is is given by pairs of score values $s_{i}$ and their probabilities $\mu[S=s_{i}]=\pi_{i}>0$, $i=1,\ldots,n$. Then it follows from the AUC-definition and properties in Section 3.3 that
$AUC_{S}\ =\ \frac{\sum_{i=2}^{n}\pi_{i}\,s_{i}\Big{(}\frac{1}{2}\,\pi_{i}(1-s_{i})+\sum_{k=1}^{i-1}\pi_{k}\,(1-s_{k})\Big{)}}{\Big{(}1-\sum_{j=1}^{n}\pi_{j}\,s_{j}\Big{)}\sum_{j=1}^{n}\pi_{j}\,s_{j}}$ (13)
@@
We use (13) to compute imlied AUC both under the source distribution for $S=\eta_{P}(X)$ with $\mu=P$ and under the target distribtion for $S=\eta_{Q}(X)=T\bigl{(}\eta_{P}(X)\bigr{)}$ with $\mu=Q$. Note that $\eta_{P}(X)$ under $P$ and $\eta_{Q}(X)$ under $Q$ are both auto-calibrated probabilistic classifiers (Proposition 1 of Vaicenavicius et al. [27]).
0.A.2 Adapting discrete distributions for QMM
@@
Consider for the real-valued score or probabilistic classifier $S$ with discrete distribution $\mu[S=s_{i}]=\pi_{i}$, $i=1,\ldots,n$, its distribution function $G(s)=\mu[S\leq s]$ for $s\in\mathbb{R}$. If the $s_{i}$ are increasingly ordered with $s_{1}<\ldots<s_{n}$, then we obtain $G(s_{i})=\sum_{j=1}^{i}\pi_{j}$, for $i=1,\ldots,n$. In particular, it follows $G(s_{n})=1$ such that $\Phi^{-1}\bigl{(}G(s_{n})\bigr{)}=\infty$ for the standard normal distribution function $\Phi$ and for the logistic distribution function.
@@
For the purpose of this paper, to avoid this problem we apply the workaround proposed by van der Burgt [4]. He suggested replacing the distribution function $G$ with the mean of itself and its left-continuous version, i.e. with $G^{\ast}$ defined by $G^{\ast}(s)=\frac{1}{2}\,\bigl{(}\mu[S\leq s]+\mu[S<s]\bigr{)}$, for $s\in\mathbb{R}$.

### v3

@@ 5 Illustration
  • The mean of the square root of the class 1 posterior probabilities as an example involving a concave function of the posterior probabilities.
For the example discrete source and target distributions of a feature (called score in the following) with values in an ordered set with 17 elements were chosen. These distributions may be interpreted as empirical distributions of samples with many ties or as the genuine distributions of discrete scores or ratings. For instance, the major credit rating agencies Standard & Poor’s, Moody’s and Fitch use rating scales with 17 to 19 different grades.
The source distribution is fully specified as follows:
  • (1)
    The conditional feature distribution for class 0 is a binomial distribution with success probability 0.4, the conditional feature distribution for class 1 is a binomial distribution with success probability 0.55. The number of trials for both binomial distributions is 16 such that the support of the distribution is the set $\{0,1,\ldots,16\}$.
  • (2)
    The class 1 prior probability is $p=0.01$.
The target distribution is incompletely specified as follows:
  • (1)
    The unconditional feature distribution is a binomial distribution with number of trials 16 whose success probability is a Vasicek-distributed random variable (Section 2.2 of Meyer, 2009) with mean 0.3 and correlation parameter 0.3.
  • (2)
    The target class 1 prior probability is $q=0.05$.
Figure 1 shows [^3] Calculations were performed with R (R Core Team, 2024). Details of some more involved calculations are described in Appendices A.1 and A.2. The R-scripts can be downloaded from https://www.researchgate.net/profile/Dirk-Tasche. the source and target unconditional score distributions. They were intentionally chosen to be quite different such that any distribution shift assumed for the recalibration must be significant.
Figure 1: Source and target score distributions for the illustrative example.
@@
Figure 2: Class 1 posterior probabilities for the illustrative example.
In Figure 2, the scale of the vertical axis is logarithmic. The dots for the probabilities have been connected with straight lines for better readability. ‘Source’ refers to the posterior probabilities in the source distribution without recalibration.
All target posterior probabilities shown in Figure 2 are well above the source posterior probabilities. This does not come as a surprise given that the target class 1 prior probability of 5% is much higher than the source class 1 prior probability of 1%. Otherwise, there are two groups of target posterior probability curves: The curves based on capped scaling, label shift and FJS on the one hand, and the curves based on the five QMM methods described in Sections 4.4 and 4.5 on the other hand. The curves of the former group are rather steep compared to the curves of the latter group.
Table 1 displays three characteristics for the source distribution as well as for the eight different target distributions that result from the recalibration methods discussed in Section 4. The row ‘Source’ shows the values for the source distribution without recalibration. The numbers in all other rows refer to the target distribution. ‘mean(probs)’ is the class 1 prior probability calculated as mean of the recalibrated posterior probabilities. ‘AUC’ is the area under the ROC curve implied by the recalibrated posterior probabilities. ‘mean(sqrt(probs))’ is the mean of the square root of the recalibrated posterior probabilities.
Table 1: Results for the application the recalibration methods to the illustrative example.
@@
Concluding from the entries in the column ‘mean(probs)’ of Table 1, only the recalibration method ‘label shift’ – introduced in Section 4.2 – is unreliable in so far as it does not achieve the required target class 1 prior probability $p=0.05$. With all seven other recalibration methods the target prior probability is matched or nearly matched.
Column ‘AUC’ of Table 1 is more varied than column ‘mean(probs)’. For the five QMM methods from Sections 4.4 and 4.5, AUC is essentially the same as the AUC of 0.802 under the source distribution. This is a consequence of the QMM design since one of the moment matching objectives is hitting the source AUC. Thanks to Proposition 4.1, one might expect also the ‘label shift’ AUC to match the source AUC. However, Proposition 4.1 is not applicable to the example of this section because by design the target feature distribution is not a mixture of the source class-conditional distributions. In any case, the high AUC values for methods ‘capped scaling’, ‘label shift’ and ‘FJS’ cause the greater slopes of their posterior probability curves in Figure 2 compared to the slopes of the QMM posterior probability curves.
As an example for the application of a concave function $C$ to the target posterior probabilities and the related expected value $E_{Q}\bigl[C(\eta_{Q}(X))\bigr]$ under the target distribution, choose the function $C(u)=\sqrt{u}$, $u\in[0,1]$. Note that (2) implies $0.05=q\leq E_{Q}\bigl[\sqrt{\eta_{Q}(X)}\bigr]\leq\sqrt{q}\approx 0.2236$.
@@ 6 Conclusions
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. This paper presents analyses and methods for recalibration from a distribution shift perspective. In order to deal with the fact that there is no unique solution for the recalibration problem, the impact of the recalibration method on the values of concave or nearly concave functions of the posterior probabilities was investigated as an additional criterion to identify meaningful solutions. It turned out that distribution shift assumptions linked to the performance of the probabilistic classifier in terms of the area under the curve (AUC) under the source distribution are useful for the design of useful recalibration methods. Two new methods called parametric CSPD (covariate shift with posterior drift) and ROC-based QMM (quasi moment matching) were proposed and were tested together with some other methods in an example setting. The outcomes of this testing exercise suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
Acknowledgement
The author would like to thank the editorial team and an anonymous reviewer for their helpful feedback which greatly improved the quality of this paper.
Appendix A Appendix
A.1 The AUC of discrete auto-calibrated probabilitistic classifiers
@@
How to calculate implied AUC for an auto-calibrated [^4] The probabilistic classifier $S$ is auto-calibrated if for all $s$ in the range of $S$ it holds that $\mu[Y=1\,|\,S=s]=s$ (Tsyplakov, 2013, Section 2.2). probabilistic classifier $S$ under a joint distribution $\mu(S,Y)$ of $S$ and $Y\in\mathcal{Y}=\{0,1\}$? Here it is assumed that the distribution of $S$ under distribution $\mu$ is discrete, i.e. it is is given by pairs of score values $s_{i}$ and their probabilities $\mu[S=s_{i}]=\pi_{i}>0$, $i=1,\ldots,n$. Then it follows from the AUC-definition and properties in Section 3.3 that
$AUC_{S}\ =\ \frac{\sum_{i=2}^{n}\pi_{i}\,s_{i}\Big(\frac{1}{2}\,\pi_{i}(1-s_{i})+\sum_{k=1}^{i-1}\pi_{k}\,(1-s_{k})\Big)}{\Big(1-\sum_{j=1}^{n}\pi_{j}\,s_{j}\Big)\sum_{j=1}^{n}\pi_{j}\,s_{j}}.$ (13)
@@
(13) is used to compute implied AUC both under the source distribution for $S=\eta_{P}(X)$ with $\mu=P$ and under the target distribution for $S=\eta_{Q}(X)=T\bigl(\eta_{P}(X)\bigr)$ with $\mu=Q$. Note that $\eta_{P}(X)$ under $P$ and $\eta_{Q}(X)$ under $Q$ are both auto-calibrated probabilistic classifiers (Proposition 1 of Vaicenavicius et al., 2019).
A.2 Adapting discrete distributions for QMM
@@
Consider for the real-valued score or probabilistic classifier $S$ with discrete distribution $\mu[S=s_{i}]=\pi_{i}$, $i=1,\ldots,n$, its distribution function $G(s)=\mu[S\leq s]$ for $s\in\mathbb{R}$. If the $s_{i}$ are increasingly ordered with $s_{1}<\ldots<s_{n}$, it follows that $G(s_{i})=\sum_{j=1}^{i}\pi_{j}$, for $i=1,\ldots,n$. In particular, it follows $G(s_{n})=1$ such that $\Phi^{-1}\bigl(G(s_{n})\bigr)=\infty$ for the standard normal distribution function $\Phi$ and for the logistic distribution function.
@@
For the purpose of this paper, to avoid this problem the workaround proposed by Van der Burgt (2008) is applied. He suggested replacing the distribution function $G$ with the mean of itself and its left-continuous version, i.e. with $G^{\ast}$ defined by $G^{\ast}(s)=\frac{1}{2}\,\bigl(\mu[S\leq s]+\mu[S<s]\bigr)$, for $s\in\mathbb{R}$.